QUESTION 23 Find the series radius of convergence. > (x - 2n n=0 0, for all x
Answer #1

Gliven series is: È (x-2) n nro Let an= (x-2)^ using Root Test; lim nad lanlyn =lim 16x-2)<1 foo convergence nico 18-2/<1 =R.